### v1

@@ Abstract
Unmanned Aerial Systems (UASs) or drones become more and more commercially available and cheap. The availability of these cheap commercial off the shelf (COTS) UASs in recent years led to potentially more threats to the traditional perimeter defense of civilian and military facilities, critical infrastructures and public events. Moreover, Law Enforcement Agencies (LEAs) find themselves confronted with the novel task of having to police the access to the lower airspace as more and more drones become available. Equally, there has been much emphasis on developing and deploying Counter-UAS systems (UASs) with Detection Tracking and Identification (DTI) solutions. However, the capabilities of these systems are hard to benchmark. Performance claims of these systems are currently not supported by evidence. In addition, no standard test methodologies are available for these DTI systems and different test methodologies make comparison of these systems hard or impossible. We report on the definition, development and verification of an objective-driven test method and corresponding comparative performance evaluation for commercial DTI solutions for C-UASs. The developed methodology is based on end-user scenarios that are operationally relevant. The test methodology is based on a generic DTI system lay-out and is detailed towards detection, tracking and identification, taking into account contextual information and end-user input. The comparative performance evaluation is developed to enable the use of the methodology in a relevant environment, thereby taking into account any potential environmental aspect that might influence DTI system performance. A comprehensive process has been outlined to develop and implement the objective-driven test method and the comparative performance evaluation. This process deals with the design of the testing system where defining the objectives the tests are addressed unambiguously. These objectives are formulated in an objective function that deals with key performance indicators for the system under test. To ensure that tests happen under comparable conditions a controlled testing environment is set up. This controlled environment covers relevant parameters and real-world aspects needed for a test scenario. Finally, such a system has to be improved and validated iteratively, meaning that once completed the test system should be applied to trials where a variety of systems are tested and compared. The work presented in this paper has firstly been verified in a simulation environment where a number relevant scenarios were used and the output the simulation injected into the testing system. Validation of the work in a relevant environment has been done in two operational trials. The operational trial results show that the method allows for performance evaluation at component level (i.e., detection, tracking or identification component) and at system level (combinations of these components and integrated DTI system of system solutions).
@@ 1 Introduction
In the Ukraine, UAVs have become a consumable resource (as opposed to a prized possession), to be used (quite literally sometimes) like an artillery shell. In this conflict, the majority of drones is currently operated by first-person view (FPV), meaning they are piloted (as opposed to autonomous). With the proliferation of such FPV-vehicles and the growing community of experienced pilots comes the reality that also our civilian airspaces are increasingly subject to such vehicles Tadić et al. (2024). It is just a question of time until unmanned assets, especially Unmanned Aerial Vehicles (UAVs) are also used for nefarious purposes European Commission and Joint Research Centre et al. (2023); Seidaliyeva et al. (2024).
As a result, the industry has reacted by advancing the field of detection, tracking and identification (DTI) systems. These systems, integral to both civilian and defense applications European Commission and Joint Research Centre et al. (2023), range from advanced radio detection and ranging (RADAR) setups to sophisticated software solutions for UAV oversight. Despite their technological advancements and increasing adoption, the evaluation of DTI systems remains a complex challenge in need of objective benchmarks, comprehensive performance metrics, and adaptable testing frameworks. With commercial interests at stake, competition may be fierce and any consultant advising a customer for one and against another system (a) better have their facts straight and (b) be prepared to defend this decision against disgruntled commercial parties.
@@
This paper introduces a comprehensive testing methodology tailored to the nuanced requirements of DTI systems. By weaving together a framework that prioritizes objective-driven test methods with the comparative analysis of system performance, our approach aims to offer a versatile and rigorous means of evaluating DTI systems. Through a meticulous integration of key performance indicators Key Performance Indicators (KPIs), weighting of these KPIs, and scenario-based testing, this methodology not only facilitates a nuanced assessment of DTI capabilities but also empowers the testers to tailor the evaluation process to their specific operational needs.
@@ 1.1 Motivation
As UASs or drones become more and more commercially available and cheap, Law Enforcement Agencies (LEAs) find themselves confronted with the novel task of having to police the access to the lower airspace. The availability of cheap commercial off the shelf (COTS) UASs in recent years Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014) has indeed led to potentially more threats European Commission and Joint Research Centre et al. (2023) to the traditional perimeter defense of civilian and military facilities, critical infrastructures and public events Tan et al. (2021). Equally, there has been much emphasis on developing and deploying Counter-UAS systems (C-UASs) with DTI solutions. However, the capabilities of these systems are hard to benchmark.
These C-UASs provide a wide range of capabilities including DTI functionalities. Performance claims of these systems are currently not supported by evidence. No standard test methodologies are available and different test methodologies make comparison of these systems hard or impossible. Each operational environment may require different DTI capabilities or configurations. This may lead to DTI systems performing differently given distinct use case scenarios and environments. The need for the LEAs to have an objective driven test methodology that enables comparative performance evaluation of DTI systems. It is of paramount importance for the LEAs to have a such a test method including performance evaluation at their disposal.
@@ 1.2 Scope
Operational end-user needs to mitigate the threat posed by UASs (i.e., drones); and
@@
Based on these elements, an integral test method including comparative performance evaluation is developed that allows for a fair comparison between different DTI systems. This test method has been verified during three user-scripted demonstration trials. In doing so, the test method fosters a better understanding among LEAs of the capabilities and performance of DTI systems in relevant operational use cases.
@@ 1.4 Paper Structure
Figure 1 provides a visual overview over the paper. We start by providing the necessary background with regard to objective driven and comparative performance testing in Section 2 and discuss building blocks of DTI systems in Section 3, which also discusses the methodology. This continues into Section 4 where we report on our approach to realize objective-driven, comparative performance evaluation for commercial DTI solutions. Section 5 touches on how we validated our approach and Section 6 closes with a conclusion and some reflections.
@@ 2.1 Introduction
When designing a performance evaluation procedure for a specific class of commercial systems, two fundamental concerns should be addressed: the process should be objective driven and the evaluation should be comparative. Both are necessary to ensure that the so evaluated systems are compared in an efficient and effective manner that is suitable to reveal whether these systems meet relevant operational needs and inform e.g., purchasing decisions towards a specific customer or end-user need.
@@ 2.1.1 Objective driven test methods
Objective-driven test methods are structured around clearly defined goals and outcomes that a product, system, or process premisses to achieve. This begins with the identification of the specific objectives the end-user has in mind, examples for this include assessing the functionality of a system, evaluating the system’s performance, verifying system reliability, or stress-testing security aspects of a system. Depending on the focus of the performance evaluation, this should then guide the subsequent design and development of test criteria and procedures as well as the definition of performance metrics. By focusing on the objectives of the evaluation process from the end-user perspective we can ensure that the outcomes align with the system’s intended purpose and the end-user requirements.
An objective driven test method should first identify the key functional requirements of the to be tested (and compared) systems and consider appropriate performance benchmarks for the application. The key functional requirements are derived from relevant operational needs. In order to test different objectives relatively independently test scenarios should be designed. By subjecting a system to these scenarios all requirements and critical features can be evaluated. In addition, the decision to use objective driven performance evaluation facilitates a transparent process that can be understood and agreed upon by developers, stake-holders and end-users.
@@ 2.1.2 Comparative performance evaluation
In comparative testing the performances of a number of test subjects or systems are compared, using a standardized set of criteria. It creates, in fact, a rating or an ordering of the tested entities depending on clearly defined operational needs; this allows for the results to be used directly for informed decisions with regard to the choice or purchase of system over the others. Comparative testing can be used to facilitate a transparent, fair and unbiased comparison of the tested entities by subjecting them to identical tests and test situations and by doing so under controlled conditions.
By designing scenarios and test cases that reflect the intended deployment of the system in the tester’s specific use-case the performance comparison does compare systems with a focus on the aspects relevant to the end-user. In many cases, the ultimate decision for a system and against another will be a trade-off, meaning that it is often not the case that a single system outperforms all others in all aspects and conditions. It is very likely that performance in one category comes with either a detrimental impact on another or with an exceedingly high cost. Often one does not need the best system at all cost but simply a system that is good enough. Comparative testing enables the tester to choose from those options which yield the highest return, meaning deliver the desired service while all relevant aspects (e.g., performance, cost) are taken into account.
@@ 2.1.3 Objective driven, comparative performance testing
Objective driven testing is particularly relevant in early stages of a project as it forces the tester to define the objectives at the outset. This inherently goal-oriented approach therefore drives clarity and facilitates transparent goals. It is introspective in that it focuses on the system’s capability to meet its intended purpose. Comparative testing on the other hand is entirely concerned with the measurable and quantifiable performance of a system in certain conditions as it will compare different systems in the basis of these metrics. While the former ascertains that a system can indeed perform the services expected from it, the latter compares the degree with which these services are performed across systems.
@@ 2.2 Important process stages
The process of designing and developing a comparative performance evaluation system can be split in three high level stages: the design of the testing system, the actual use of the system (e.g., by means of an opeational trial) and, ultimately, the validation of the system; cf. Figure 2.
With regard to the design of such a system or process, we already discussed above the need for clearly stated objectives to unambiguously define what the goals of the testing process are. With that in place, we can discuss the metrics we want to use. These metrics must cover all aspects the tester (user) considers relevant for their eventual choice. Examples for such metrics are e.g., system performance and reliability, used hardware, response time of sensors, ease of installation, ease of use or cost (both OPEX as well as CAPEX).
To ensure that tests happen under comparable conditions a controlled testing environment needs to be set up. This controlled environment should have all relevant parameters and real-world aspects needed for a test scenario (of which there are a number) covered. With regard to the testing scenarios, these should be designed to be generic and abstract enough to cover entire classes of groups of real-world scenarios but have to be specific enough to collectively cover the breadth of scenarios, including special cases and outliers. Note that scenarios should be designed to be tunable to specific needs and interests. For example, a scenario may revolve around a train station, in which case the scenario should be an abstract representation of such a building. It should be tunable to a small station during peak traffic as well as to a major station in the middle of the night.
With the above in place, a number of commercially competing products can be subjected to tests and their resulting performances can be compared. Note that preferences for specific metrics over others (maybe reliability is more important than cost-effectiveness) are expressed by the tester and are to be reflected in an objective function which is applied to all partial results for each tested system.
Finally, such a system has to be improved and validated iteratively, meaning that once completed the system should be applied to trials where a variety of systems are tested and compared. Emphasis should be placed in the correctness of the outcome as well as the transparency of the process as well as the easy of use.
@@ 2.3 System Design Principles
In order to be broadly applicable, the evaluation procedure must be comprehensive in that it is capable of evaluation a range of performance metrics and can so over a number of diverse scenarios. This should cover the application domain and spectrum of tested systems and represent all major use-cases (in abstract scenarios). The evaluation should be impartial and void of inherent biases (towards specific manufacturers, systems, paradigms or users) and should therefore be able to use any metric and run any scenario when testing a system; metrics and scenarios may not favor any individual system under investigation.
@@
Finally, the evaluation system has to be available (both in access to the system as well as in ease of use thereof) to the broader community interested in the evaluation of a specific system. If access to the performance comparison is restricted to certain stake-holders then the main goal (that of providing a objective means to fairly assess the performance of systems to a specific use case) cannot be accomplished.
@@ 2.4 Defining the objective
An objective function in the context of evaluating complex systems like DTI systems quantitatively represents the user’s priorities and performance aspects they wish to focus on. This objective function’s result is a weighted sum of normalized performance metrics, where each weight corresponds to the user’s emphasis (preference?) for that metric. Users begin by identifying KPIs relevant to their operational needs, such as detection and track accuracy, and identification reliability. Each metric is then assigned a weight based on its importance to the user’s specific application or mission requirements. The evaluation system provides a user-friendly interface for this process, offering templates for common use cases while allowing custom configurations. The objective function is designed to be flexible, accommodating changes in user priorities and the introduction of new KPIs as DTI technologies evolve.
The overall aim is to design an evaluation system that can be tailored by a user to compare commercial systems for a broad set of specific use-cases and scenarios. A user must therefore be able to tune the scenario to represent their use case and must further be able to change the scoring of the system to reflect the user’s priorities and preferences. The latter is captured in the objective function of the performance evaluation, which quantitatively represents the user’s priorities and focuses on those performance aspects the user wishes to focus on.
The objective function commonly uses normalized test scores for specific metrics, weighing them according to the user’s preferences. This function should be configurable for the user, and the system should facilitate this in an easy and intuitively understandable manner.
@@ 3.1 Operational needs and scenario modelling
Relevant operational needs and requirements form the basis in defining functional requirements the DTI systems should meet. The key underlying operational requirements can be summarized as the need to detect and counter any drone, of any size, exhibiting certain behaviour in different environmental conditions. Scenario modelling for evaluating complex systems like DTI involves the creation of a comprehensive set of environments and conditions under which the systems’ performances are assessed. These scenarios are designed to mimic a broad spectrum of real-world operational conditions, including various climates, terrains UAS malicious and neutral behaviour, and urban density levels. Each scenario is characterized by specific parameters such as weather conditions, the presence of obstacles, and the density of non-target entities (e.g., birds or non-malicious drones) to assess systems’ adaptability and reliability under varied conditions. The process incorporates input from industry experts, operational data, and historical incident reports to ensure relevance and comprehensiveness. This approach ensures that the evaluation system can test DTI systems’ performance in controlled yet realistic and relevant conditions.

### v2

@@ Abstract
Unmanned Aerial Systems (UASs) or drones become more and more commercially available and cheap. The availability of these commercial off the shelf (COTS) UASs in recent years led to potentially more threats to the traditional perimeter defense of civilian and military facilities, critical infrastructures and public events. Moreover, Law Enforcement Agencies (LEAs) find themselves confronted with the novel task of having to police the access to the lower airspace as more and more drones become available. Equally, there has been much emphasis on developing and deploying Counter-UAS systems (C-UASs) with detection, tracking and identification (DTI) solutions. However, the capabilities of these systems are hard to benchmark. Performance claims of these systems are currently not supported by evidence. In addition, no standard test methodologies are available for these DTI systems and different test methodologies make comparison of these systems hard or impossible. We report on the definition, development and verification of an objective-driven test method and corresponding comparative performance evaluation for commercial DTI solutions for C-UASs. The developed methodology is based on end-user scenarios that are operationally relevant. The test methodology is based on a generic DTI system lay-out and is detailed towards detection, tracking and identification, taking into account contextual information and end-user input. The comparative performance evaluation is developed to enable the use of the methodology in a relevant environment, thereby taking into account any potential environmental aspect that might influence DTI system performance. A comprehensive process has been outlined to develop and implement the objective-driven test method and the comparative performance evaluation. This process deals with the design of the testing system where defining the objectives the tests are addressed unambiguously. These objectives are formulated in an objective function that deals with key performance indicators for the system under test. To ensure that tests happen under comparable conditions a controlled testing environment in terms of use cases and UAS behaviour is set up. This controlled environment covers relevant parameters and real-world aspects needed for a test scenario. Finally, such a system has to be improved and validated iteratively, meaning that once completed the test system should be applied to trials where a variety of systems are tested and compared. The work presented in this paper has firstly been verified in a simulation environment where a number relevant scenarios were used and the output the simulation injected into the testing system. Validation of the work in a relevant environment has been done in two operational trials. The operational trial results show that the method allows for performance evaluation at component level (i.e., detection, tracking or identification component) and at system level (combinations of these components and integrated DTI system of system solutions).
@@ 1 Introduction
In the Ukraine, UASs have become a consumable resource (as opposed to a prized possession), to be used (quite literally sometimes) like an artillery shell. In the current Ukraine conflict, the majority of drones is currently operated by first-person view (FPV), meaning they are piloted (as opposed to being autonomous). With the proliferation of such FPV-vehicles and the growing community of experienced pilots comes the reality that also our civilian airspaces are increasingly subject to such vehicles Tadić et al. (2024). It is just a question of time until unmanned assets, especially UASs are also used for nefarious purposes European Commission and Joint Research Centre et al. (2023); Seidaliyeva et al. (2024).
In response, the industry has steadily advanced the field of detection, tracking and identification (DTI) systems. These systems, integral to both civilian and defense applications European Commission and Joint Research Centre et al. (2023), range from advanced RADAR setups to sophisticated software solutions for Unmanned Aerial Vehicle (UAV) oversight. Despite their technological advancements and increasing adoption, the evaluation of DTI systems remains a complex challenge in need of objective benchmarks, comprehensive performance metrics, and adaptable testing frameworks. With commercial interests at stake, competition may be fierce and any consultant advising a customer for one and against another system (a) better have their facts straight and (b) be prepared to defend this decision against disgruntled commercial parties.
@@
This paper introduces a comprehensive testing methodology tailored to the nuanced requirements of DTI systems. By weaving together a framework that prioritizes objective-driven test methods with the comparative analysis of system performance, our approach aims to offer a versatile and rigorous means of evaluating DTI systems. Through a meticulous integration of Key Performance Indicators (KPIs), the subsequent weighting of these KPIs, and the use of scenario-based testing, this methodology not only facilitates a nuanced assessment of DTI capabilities but also empowers the testers to tailor the evaluation process to their specific operational needs.
@@ 1.1 Motivation
As UASs, or drones, become more and more commercially available and cheap, LEAs find themselves confronted with the novel task of having to police the access to the lower airspace. The availability of cheap commercial off the shelf (COTS) UASs in recent years Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014) has indeed led to potentially more threats European Commission and Joint Research Centre et al. (2023) to the traditional perimeter defense of civilian and military facilities, critical infrastructures and public events Tan et al. (2021). Equally, there has been much emphasis on developing and deploying C-UASs with DTI solutions. However, the capabilities of these systems are hard to benchmark.
These C-UASs provide a wide range of capabilities including DTI functionalities. Performance claims of these systems are currently not supported by evidence. No standard test methodologies are available and different test methodologies make comparison of these systems hard or impossible. Each operational environment may require different DTI capabilities or configurations. This may lead to DTI systems performing differently given distinct use case scenarios and environments. There is an urgent need for the LEAs to have an objective driven test methodology that enables comparative performance evaluation of DTI systems. It is of paramount importance for the LEAs to have a such a test method including performance evaluation at their disposal.
@@ 1.2 Scope
Operational end-user needs for specific mitigation techniques to counter UASs; and
@@
Based on these elements, an integral test method including comparative performance evaluation is developed that allows for a fair comparison between different DTI systems. This test method has been verified during three user-scripted demonstration trials. In doing so, the test method fosters a better understanding among LEAs of the capabilities and performance of DTI systems in relevant operational use cases. In addition, relevant feedback from theLEAs can be integrated into the test methodology.
@@ 1.4 Paper Structure
Figure 1 provides a visual overview over the paper. We start by providing the necessary background with regard to objective driven and comparative performance testing in Section 2. We discuss building blocks of DTI systems and the methodology in Section 3. This continues into Section 4 where we report on our approach to realize objective-driven, comparative performance evaluation for commercial DTI solutions in detail. Section 5 touches on how we validated our approach and Section 6 closes with a conclusion and some reflections. Formal details about the metrics used are provided in the appendices.
@@ 2.1 Introduction
When designing a performance evaluation procedure for a specific class of commercial systems, two fundamental concerns should be addressed: the process should be objective driven and the evaluation should be comparative. Both are necessary to ensure that the so evaluated systems are compared in an efficient and effective manner that is suitable to reveal whether these systems meet relevant operational needs and to make informed (e.g., purchasing) decisions towards a specific customer or end-user need.
@@ 2.1.1 Objective driven test methods
Objective-driven test methods are structured around clearly defined goals and outcomes that a product, system, or process promises to achieve. This begins with the identification of the specific objectives the end-user has in mind, examples for this include assessing the functionality of a system, evaluating the system’s performance, verifying system reliability, or stress-testing security aspects of a system. Depending on the focus of the performance evaluation, this should then guide the subsequent design and development of test criteria and procedures as well as the definition of performance metrics. By focusing on the objectives of the evaluation process from the end-user perspective we can ensure that the outcomes align with the system’s intended purpose and the end-user requirements.
An objective driven test method should first identify the key functional requirements of the systems to be tested (and compared). It should also consider appropriate performance benchmarks for the application. The key functional requirements are derived from relevant operational needs. In order to test different objectives relatively independently test scenarios should be designed. By subjecting a system to these scenarios all requirements and critical features can be evaluated. In addition, the decision to use objective driven performance evaluation facilitates a transparent process that can be understood and agreed upon by developers, stakeholders and end-users.
@@ 2.1.2 Comparative performance evaluation
In comparative testing the performances of a number of test subjects or systems are compared, using a standardized set of criteria. It creates, in fact, a rating or an ordering of the tested entities depending on clearly defined operational needs; this allows for the results to be used directly for informed decisions with regard to the choice or purchase of a system over the others. Comparative testing can be used to facilitate a transparent, fair and unbiased comparison of the tested entities by subjecting them to identical tests and test situations and by doing so under controlled conditions.
By designing scenarios and test cases that reflect the intended deployment of the system in the tester’s specific use-case the performance comparison does compare systems with a focus on the aspects relevant to the end-user. In many cases, the ultimate decision for a system against another will be a trade-off, meaning that it is often not the case that a single system outperforms all others in all aspects and conditions. It is very likely that performance in one category comes with either a detrimental impact on another or with an excessive high cost. Often one does not need the best system at all cost but simply a system that meets the user requirements. Comparative testing enables the tester to choose from those options which yield the highest return, meaning deliver the desired service while all relevant aspects (e.g., performance, cost) are taken into account.
@@ 2.1.3 Objective driven, comparative performance testing
Objective driven testing is particularly relevant in early stages of a project as it forces the tester to define the objectives of a specific test at the outset. This inherently goal-oriented approach therefore drives clarity and facilitates transparent goals. It is introspective in that it focuses on the system’s capability to meet its intended purpose. Comparative testing on the other hand is entirely concerned with the measurable and quantifiable performance of a system in certain conditions as it will compare different systems in the basis of these metrics. While the former ascertains that a system can indeed perform the services expected from it, the latter compares the degree with which these services are performed across systems.
@@ 2.2 Important process stages
Designing and developing a comparative performance evaluation system happens over three high level stages: the design of the testing system, the actual use of the system (e.g., by means of an operational trial) and the validation of the system; cf. Figure 2.
With regard to the design of such a system or process, we already discussed above the need for clearly stated objectives to unambiguously define what the goals of the testing process are. With that in place, we can discuss the metrics we want to use. These metrics must cover all aspects the tester (user) considers relevant for their eventual choice. Examples for such metrics are e.g., system performance and reliability, used hardware, response time of sensors, ease of installation, ease of use or cost (both, OPEX and CAPEX).
To ensure that tests happen under comparable conditions a controlled testing environment needs to be set up. This controlled environment should have all relevant parameters and real-world aspects needed for a test scenario (of which there are a number) covered. The testing scenarios should be generic and abstract as well as specific enough to collectively cover the breadth of real-world scenarios. Note that scenarios should be designed to be tunable to specific needs and interests. For example, a scenario may focus on a train station, in and contain an abstract representation of such a building. It should be tunable to a small station during peak traffic as well as to a major station in the middle of the night.
With the above in place, a number of commercially competing products can be subjected to tests and their resulting performances can be compared. Note that preferences for specific metrics over others (maybe reliability is more important than cost-effectiveness) are expressed by the end-user and are to be reflected in an objective function which is applied to all results for each tested system.
Finally, such a comparative test and evaluation methodology has to be improved and validated iteratively, meaning that once completed the system should be applied to trials where a variety of systems are tested and compared. Emphasis should be placed on the correctness of the outcome and the transparency of the process as well as the easy of use.
@@ 2.3 System Design Principles
In order to be broadly applicable, the evaluation procedure must be comprehensive in that it is capable of evaluating a range of performance metrics and can so over a number of diverse scenarios. This should cover the application domain and spectrum of tested systems and represent all major use-cases (in abstract scenarios). The evaluation should be impartial and void of inherent biases (towards specific manufacturers, systems, paradigms or users) and should therefore be able to use any metric and run any scenario when testing a system; metrics and scenarios may not favor any individual system under investigation.
@@
Finally, the evaluation system has to be available (both in access to the system as well as in ease of use thereof) to the broader community interested in the evaluation of a specific system. If access to the performance comparison is restricted to certain stakeholders then the main goal (that of providing a objective means to fairly assess the performance of systems to a specific use case) cannot be accomplished.
@@ 2.4 Defining the objective
An objective function in the context of evaluating complex systems like DTI systems quantitatively represents the user’s priorities and performance aspects they wish to focus on. This objective function’s result is a weighted sum of normalized performance metrics, where each weight corresponds to the user’s emphasis for that metric.
Users begin by identifying KPIs relevant to their operational needs, such as detection and track accuracy, and identification reliability. Each metric is then assigned a weight based on its importance to the user’s specific application or mission requirements. The evaluation system provides a user-friendly interface for this process, offering templates for common use cases while allowing custom configurations. The objective function is designed to be flexible, accommodating changes in user priorities and the introduction of new KPIs as DTI technologies evolve.
The overall aim is to design an evaluation system that can be tailored by a user to compare commercial systems for a set of specific use-cases and scenarios. The user must be able to tune the scenario to represent their use case and must further be able to change the scoring of the system to reflect their priorities and preferences. The latter is captured in the objective function of the performance evaluation, which quantitatively represents the user’s priorities and focuses on those performance aspects the user prioritizes.
The objective function commonly uses normalized test scores for specific metrics, weighing them according to the user’s preferences. This function should be configurable for the user, and the system should facilitate this in an easy and intuitive manner.
@@ 3.1 Operational needs and scenario modelling
Relevant operational needs and requirements form the basis in defining functional requirements the DTI systems should meet. The key underlying operational requirements can be summarized as the need to detect and counter any drone, of any size, exhibiting certain behaviour in different environmental conditions. Note that the countering UASs is not in the scope of this work.
Scenario modelling for evaluating complex systems like DTI involves the creation of a comprehensive set of environments and conditions under which the systems’ performances are assessed. These scenarios are designed to mimic a broad spectrum of real-world operational conditions, including various climates, terrains, UAS malicious and neutral behaviour, and urban density levels. Each scenario is characterized by specific parameters such as weather conditions, the presence of obstacles, and the density of non-target entities (e.g., birds or non-malicious drones) to assess systems’ adaptability and reliability under varied conditions. The process incorporates input from industry experts, operational data, and historical incident reports to ensure relevance and comprehensiveness. This approach ensures that the evaluation system can test DTI systems’ performance in controlled yet realistic and relevant conditions.